# Multi-object Data Integration in the Study of Primary Progressive Aphasia

### v1

@@ 3.2 Prior Distribution on Multi-Modal Coefficients
Our integrated prior construction on coefficients ${\boldsymbol{\beta}}_{p}$’s and $\{\theta_{p,p^{\prime}}:p<p^{\prime}\}$ for multi-modal predictors is fundamental to exploiting topology of the image objects and cross-information among them by forming principled linkages among images at the ROI level. The prior construction is aimed at: (a) identification of influential ROIs with uncertainty; (b) shrinkage of unimportant voxel coefficients to zero within an influential ROI; and (c) guaranteeing efficient computation of the posterior for the proposed prior. We cast the problem of identifying influential ROIs from the multi-modal images as a high-dimensional variable selection problem and formulate prior distributions on multi-modal object coefficients building upon the existing literature on prior constructions for high-dimensional regression coefficients.
To this end, two classes of prior distributions on coefficients are typically employed in high-dimensional regression literature. The traditional approach is to develop a discrete mixture of prior distributions (George and McCulloch, 1993; Scott and Berger, 2010). These methods enjoy the advantage of inducing exact sparsity for a subset of parameters but may face computational challenges when the number of coefficients is large. As an alternative to this approach, a continuous approximation to the discrete mixture priors (Carvalho et al., 2010) have emerged which induce approximate sparsity in high-dimensional parameters. Such prior distributions can mostly be expressed as global-local scale mixtures of Gaussians (Polson and Scott, 2010), are computationally efficient, and offer an approximation to the operating characteristics of discrete mixture priors.
@@
The direct application of variable selection prior to multi-modal coefficients is unappealing for multiple reasons. First, an ordinary variable selection prior on coefficients ${\boldsymbol{\Theta}}$ and ${\boldsymbol{\beta}}_{p}$’s identifies cells in ${\boldsymbol{A}}_{i}$ and ${\boldsymbol{g}}_{i,p}$ (which in our application refers to the network edges and GM voxels) significantly related to the predictor, rather than identifying influential ROIs. Second, we seek to impose two additional restrictions on the prior construction of ${\boldsymbol{\Theta}}$ and ${\boldsymbol{\beta}}_{p}$ motivated by the neuro-scientific application. First, if at least one of the $p$th and $p^{\prime}$th ROIs is not related to the speech rate, the edge coefficient $\theta_{p,p^{\prime}}$ corresponding to the edge between $p$th and $p^{\prime}$th nodes is unimportant. Second, if the $p$ ROI is deemed uninfluential then all voxels within the $p$th ROI are unrelated to the speech rate, i.e., ${\boldsymbol{\beta}}_{p}={\boldsymbol{0}}$. These restrictions are relevant due to the hierarchical arrangement of voxels and ROIs and are jointly referred to as the hierarchical constraint. Finally, we expect the matrix of coefficients ${\boldsymbol{\Theta}}$ (which itself can be regarded as describing a weighted network) to exhibit transitivity effects, that is, we expect that if the interactions between regions $p$ and $p^{\prime}$ and between regions $p^{\prime}$ and $p^{\prime\prime}$ are both related to the speech rate, the interaction between regions $p$ and $p^{\prime\prime}$ will likely be associated with the speech rate. The property of transitivity reflects the assumption of small-world network structure which is a popular model for neural organization that achieves the simultaneous demands of network integration and segregation (Bassett and Bullmore, 2006) and has been considered in prior studies of PPA (Nigro et al., 2022; Tao et al., 2020). An ordinary variable selection prior on multi-modal coefficients does not necessarily conform to all these requirements.
@@
where $\delta_{0}$ corresponds to the Dirac-delta function, $\lambda_{p,p^{\prime}}$ is a local shrinking parameter corresponding to the $(p,p^{\prime})$th edge and $\sigma_{\theta}$ is the global shrinking parameter for the network coefficient. The prior closely mimics the spike-and-slab variable selection structure with an important difference. While an ordinary spike-and-slab prior introduces a binary inclusion indicator corresponding to each variable, (4) enforces $\theta_{p,p^{\prime}}=0$ when either $\xi_{p}=0$ or $\xi_{p^{\prime}}=0$. Such a formulation is sensible from a network perspective as it implies that the edge connecting two network nodes is unrelated to the predictor when at least one of the network nodes is not influential, i.e., it satisfies hierarchical constraint. Additionally, the formulation naturally incorporates transitivity effects in the network coefficient ${\boldsymbol{\Theta}}$. Further, the use of a single binary inclusion indicator to simultaneously determine the influence of ROIs with respect to both structural and network information is informed by prior evidence that structural atrophy and disruptions of functional connectivity can be co-localized in PPA disease progression (Brown et al., 2019) as well as other progressive NDs (Raj et al., 2015). We further assign half-Cauchy distributions on $\sigma_{\theta}\sim C^{+}(0,1)$ and $\lambda_{p,p^{\prime}}\stackrel{{\scriptstyle ind.}}{{\sim}}C^{+}(0,1)$ to complete prior specification on network coefficient. Integrating out $\sigma_{\theta}$ and $\lambda_{p,p^{\prime}}$ in (4), $\theta_{p,p^{\prime}}|\tau^{2}_{\theta},\xi_{p}=1,\xi_{p^{\prime}}=1$ follows the popular horseshoe prior (Carvalho et al., 2010) which offers a flexible prior structure for precise estimation of nonzero network edge coefficients a posteriori.
@@
for $v=1,...,V_{p};\>p=1,...,P$. The prior structure (5) induces approximate sparsity in voxel-level GM coefficients $\gamma_{v,p}$ by shrinking the components which are less influential toward zero while retaining the true signals (Polson and Scott, 2010). Finally, the binary inclusion indicators are assigned Bernoulli prior distribution $\xi_{p}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}Ber(\nu)$ with $\nu\sim Beta(a_{\nu},b_{\nu})$ to account for multiplicity correction (Scott and Berger, 2010). Notably, an estimate of the posterior probability of the event $\{\xi_{p}=1\}$ shows the uncertainty in identifying the $p$th ROI to be influential. Thus, $P(\xi_{p}=1|\mbox{Data})$ close to $1$ or 00 signifies strong evidence in favor of identifying the $p$th ROI to be active or inactive, respectively. The prior specification on multi-modal coefficients is completed by assigning an inverse Gamma $IG(a_{\tau_{\theta}},b_{\tau_{\theta}})$ and $IG(a_{\tau_{\beta}},b_{\tau_{\beta}})$ prior on the error variances $\tau^{2}_{\theta}$ and $\tau^{2}_{\beta}$ respectively. We assign $N(0,1)$ prior distribution on $\psi_{p,h}^{a},\psi_{p,h}^{g}$ for all $p=1,...,P$ and $h=1,...,H$.
Notably, the integrated prior distribution ensures significant feedback from voxel-level inference to ROI-level inference. An ROI is determined to be influential by considering the relevance of all the voxels within that ROI, along with the relevant edges of the network object. This means that multiple weak voxel-level signals within an ROI can collectively identify it as influential, or a strong signal at a single voxel can make the ROI influential. Our method achieves a principled Bayesian balance, weighing the strength of all signals contributing to that ROI. We study these aspects empirically in Section 5. Additionally, detailed inference on $\gamma_{v,p}$ enables voxel-level association between the image and the speech rate measure from our method for influential ROIs. However, we emphasize that the $\gamma_{v,p}$ values are assigned a continuous shrinkage prior, meaning they will not be exactly zero in any MCMC iterations. To determine whether a voxel is significantly related to the cognitive variable, we need to apply post-processing on the MCMC samples of $\gamma_{v,p}$ following the approach described in the literature (Li and Pati, 2017; Guha and Rodriguez, 2021). While we present inference on $\beta_{v,p}$, we do not emphasize voxel identification in the simulation study following the post-processing steps with $\gamma_{v,p}$ due to space constraint, since our main focus is identifying influential ROIs.
@@ 4 Posterior Computation
Full conditional distributions for all the parameters are available and mostly correspond to standard families (available in the supplementary file). Thus, posterior computation can proceed through a Markov chain Monte Carlo (MCMC) algorithm. The MCMC sampler is run for $5000$ iterations, with the first $1000$ discarded as burn-in. All simulation scenarios show an average effective sample size of over 3143, indicating fairly uncorrelated post-burn-in MCMC samples.
We have implemented our code in R (without using any C++, Fortran, or Python interface) on a cluster computing environment with three interactive analysis servers, 56 cores each with the Dell PE R820: 4x Intel Xeon Sandy Bridge E5-4640 processor, 16GB RAM, and 1TB SATA hard drive. Different replications of the model are implemented under a parallel architecture by making use of the packages doparallel and foreach within R.
@@ 5.1 Data Generation
For each sparsity level and co-localization degree, the coefficient corresponding to the edge connecting the $p$-th and $p^{\prime}$-th region is drawn from the following mixture distribution,
$\theta_{(p,p^{\prime}),t}|\xi^{\theta}_{p,t},\xi^{\theta}_{p^{\prime},t}\sim\xi_{p,t}^{\theta}\xi_{p^{\prime},t}^{\theta}N(\mu_{\theta,(p,p^{\prime})},\sigma_{\theta}^{2})+(1-\xi_{p,t}^{\theta}\xi_{p^{\prime},t}^{\theta})\delta_{0},\>\>\>\theta_{(p,p^{\prime}),t}=\theta_{(p^{\prime},p),t};.$ (6)
@@
(6) ensures that any edge connecting to the $p$-th region in the network response is unrelated to the predictor if the $p$-th region is un-influential, i.e., $\xi^{\theta}_{p,t}=0\Rightarrow\theta_{(p,p^{\prime}),t}=0$ for all $p^{\prime}\in\{1,..,P\}$. Similarly, corresponding to each un-influential region $\mathcal{R}_{p}$, i.e., when $\xi^{\beta}_{p,t}=0$, the $V\times 1$ dimensional GM coefficient ${\boldsymbol{\beta}}_{p,t}$ is set at ${\boldsymbol{0}}$. When $\xi^{\beta}_{p,t}=1$, i.e., the $p$-th region is influential, we randomly choose $\upsilon_{t}$ proportion of cell coefficients in the $p$-th region to be nonzero and the rest are set at zero. Thus $\upsilon_{t}$ is the density of activated voxels within a region. The nonzero coefficients within ${\boldsymbol{\beta}}_{p,t}$ are simulated from $N(\mu_{\beta,(v,p)},\sigma_{\beta}^{2})$, where the values of $\mu_{\theta,(p,p^{\prime})}$ and $\mu_{\beta,(v,p)}$ are drawn from $Unif(0.25,1)$. We fix $\sigma_{\theta}^{2}=1$, $\sigma_{\beta}^{2}=1$. We control the signal-to-noise ratio by changing the error variances $\tau_{t,\theta}^{2}$ and $\tau_{t,\beta}^{2}$.
@@
(A2) Spatial simulations: Simulating true coefficients ${\boldsymbol{\Theta}}_{t}$ and ${\boldsymbol{\beta}}_{p,t}$. We also consider simulation scenarios where structural coefficients are spatially indexed, i.e., $\beta_{v,p}=\beta({\boldsymbol{s}}_{v,p})$, with ${\boldsymbol{s}}_{v,p}$ representing the spatial location of the $v$th voxel in the $\mathcal{R}_{p}$. This simulation setting assumes perfect co-localization, i.e., $\xi_{p,t}=\xi_{p,t}^{\theta}=\xi_{p,t}^{\beta}$. The network edge coefficient $\theta_{(p,p^{\prime}),t}$ is simulated following (6). When $\xi_{p,t}=1$, we randomly choose $\upsilon_{t}$ proportion of voxel coefficients in $\mathcal{R}_{p}$ to be nonzero, and the rest are set at zero. the nonzero coefficient $\beta({\boldsymbol{s}}_{v,p})$ is simulated from a Gaussian process with mean 1 and a covariance function given by the Matérn Kernel: $cor(\beta({\boldsymbol{s}}_{v,p}),\beta({\boldsymbol{s}}_{v^{\prime},p}))=K(d|\rho,\nu)=\frac{2^{1-\nu}}{\Gamma(\nu)}\left((2\nu)^{1/2}\frac{d}{\rho}\right)^{\nu}K_{\nu}\left((2\nu)^{1/2}\frac{d}{\rho}\right),$ where $d=||{\boldsymbol{s}}_{v,p}-{\boldsymbol{s}}_{v^{\prime},p}||$ and $K_{\nu}$ is the modified Bessel function of order $\nu$. We set $\nu=2$ to ensure a fairly smooth spatial surface and $\rho=1$ for a reasonable range of spatial dependence.
@@
We simulate data under two different node sparsity parameters, $(1-\nu_{t})=0.85,0.70$. Finally, a different number of ROIs and different numbers of voxels within an ROI are considered. Specifically, we consider two cases, (a) $P=20$ and $V=10$, and (b) $P=100$ and $V=50$. Cases (a) and (b) are referred to as the “small dimensional example” and “high dimensional example.” In both these cases we have approximately similar number of parameters to estimate from the network and the structural coefficients. We also vary the signal-to-noise ratio at values $0.25,0.5,1,2,4$. All simulation scenarios are summarized in Table 1. Notably, the fitted model in Section 3 assumes co-localization. Hence, simulations under nonzero $c_{t}$ and spatial correlation between GM coefficients will allow us to assess the performance of BMRR under more realistic scenarios with model mis-specifications.

### v2

@@ 3.2 Prior Distribution on Multi-Modal Coefficients
Our integrated prior construction on coefficients ${\boldsymbol{\beta}}_{p}$’s and $\{\theta_{p,p^{\prime}}:p<p^{\prime}\}$ for multi-modal predictors is fundamental to exploiting topology of the image objects and cross-information among them by forming principled linkages among images at the ROI level. The prior construction is aimed at: (a) identification of influential ROIs with uncertainty; (b) shrinkage of unimportant voxel coefficients to zero within an influential ROI; and (c) guaranteeing efficient computation of the posterior for the proposed prior. We cast the problem of identifying influential ROIs from the images as a high-dimensional variable selection problem and formulate prior distributions on multi-modal object coefficients building upon the existing literature on prior constructions for high-dimensional regression coefficients.
@@
The direct application of variable selection prior (George and McCulloch, 1993; Carvalho et al., 2010) to multi-modal coefficients is unappealing for multiple reasons. First, an ordinary variable selection prior on coefficients ${\boldsymbol{\Theta}}$ and ${\boldsymbol{\beta}}_{p}$’s identifies cells in ${\boldsymbol{A}}_{i}$ and ${\boldsymbol{g}}_{i,p}$ (network edges and GM voxels) significantly related to the predictor, rather than identifying influential ROIs. Second, we seek to impose two additional restrictions on the prior construction of ${\boldsymbol{\Theta}}$ and ${\boldsymbol{\beta}}_{p}$ motivated by the neuro-scientific application. First, if at least one of the $p$th and $p^{\prime}$th ROIs is not related to the speech rate, the edge coefficient $\theta_{p,p^{\prime}}$ corresponding to the edge between $p$th and $p^{\prime}$th nodes is unimportant. Second, if the $p$th ROI is deemed uninfluential then all voxels within the $p$th ROI are unrelated to the speech rate, i.e., ${\boldsymbol{\beta}}_{p}={\boldsymbol{0}}$. These restrictions are relevant due to the hierarchical arrangement of voxels and ROIs and are jointly referred to as the hierarchical constraint. Finally, we expect the matrix of coefficients ${\boldsymbol{\Theta}}$ (which itself can be regarded as describing a weighted network) to exhibit transitivity effects, that is, we expect that if the interactions between regions $p$ and $p^{\prime}$ and between regions $p^{\prime}$ and $p^{\prime\prime}$ are both related to the speech rate, the interaction between regions $p$ and $p^{\prime\prime}$ will likely be associated with the speech rate. The property of transitivity reflects the assumption of small-world network structure which is a popular model for neural organization that achieves the simultaneous demands of network integration and segregation (Bassett and Bullmore, 2006) and has been considered in prior studies of PPA (Nigro et al., 2022; Tao et al., 2020). An ordinary variable selection prior on multi-modal coefficients does not necessarily conform to all these requirements.
@@
where $\delta_{0}$ corresponds to the Dirac-delta function, $\lambda_{p,p^{\prime}}$ is a local shrinking parameter corresponding to the $(p,p^{\prime})$th edge and $\sigma_{\theta}$ is the global shrinking parameter for the network coefficient. The prior closely mimics the spike-and-slab variable selection structure with an important difference. While an ordinary spike-and-slab prior introduces a binary inclusion indicator corresponding to each variable, (4) enforces $\theta_{p,p^{\prime}}=0$ when either $\xi_{p}=0$ or $\xi_{p^{\prime}}=0$. Such a formulation is sensible from a network perspective as it implies that the edge connecting two network nodes is unrelated to the predictor when at least one of the network nodes is not influential, i.e., it satisfies hierarchical constraint. Additionally, the formulation naturally incorporates transitivity effects in the network coefficient ${\boldsymbol{\Theta}}$. Further, the use of a single binary inclusion indicator to simultaneously determine the influence of ROIs with respect to both structural and network information is informed by prior evidence that structural atrophy and disruptions of functional connectivity can be co-localized in PPA disease progression (Battistella et al., 2019) as well as other progressive NDs (Johnson et al., 2000) (see Section 2 for details). We further assign half-Cauchy distributions on $\sigma_{\theta}\sim C^{+}(0,1)$ and $\lambda_{p,p^{\prime}}\stackrel{{\scriptstyle ind.}}{{\sim}}C^{+}(0,1)$ to complete prior specification on network coefficient. Integrating out $\sigma_{\theta}$ and $\lambda_{p,p^{\prime}}$ in (4), $\theta_{p,p^{\prime}}|\tau^{2}_{\theta},\xi_{p}=1,\xi_{p^{\prime}}=1$ follows the popular horseshoe prior (Carvalho et al., 2010) which offers a flexible prior structure for precise estimation of nonzero network edge coefficients a posteriori.
@@
for $v=1,...,V_{p};\>p=1,...,P$. The prior structure (5) induces approximate sparsity in voxel-level GM coefficients $\gamma_{v,p}$ by shrinking the components which are less influential toward zero while retaining the true signals (Polson and Scott, 2010). Finally, the binary inclusion indicators are assigned Bernoulli prior distribution $\xi_{p}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}Ber(\nu)$ with $\nu\sim Beta(a_{\nu},b_{\nu})$ to account for multiplicity correction. Notably, an estimate of the posterior probability of the event $\{\xi_{p}=1\}$ shows the uncertainty in identifying the $p$th ROI to be influential. Thus, $P(\xi_{p}=1|\mbox{Data})$ close to $1$ or 00 signifies strong evidence in favor of identifying the $p$th ROI to be active or inactive, respectively. The prior specification on multi-modal coefficients is completed by assigning an inverse Gamma $IG(a_{\tau_{\theta}},b_{\tau_{\theta}})$ and $IG(a_{\tau_{\beta}},b_{\tau_{\beta}})$ prior on the error variances $\tau^{2}_{\theta}$ and $\tau^{2}_{\beta}$ respectively. We assign a flat prior on $\psi_{p,h}^{a},\psi_{p,h}^{g}$ for all $p=1,...,P$ and $h=1,...,H$.
Notably, the integrated prior distribution ensures significant feedback from voxel-level inference to ROI-level inference. An ROI is determined to be influential by considering the relevance of all the voxels within that ROI, along with the relevant edges of the network object. This means that multiple weak voxel-level signals within an ROI can collectively identify it as influential, or a strong signal at a single voxel can make the ROI influential. Our method achieves a principled Bayesian balance, weighing the strength of all signals contributing to that ROI. We study these aspects empirically in Section 5. Additionally, detailed inference on $\gamma_{v,p}$ enables voxel-level association between the image and the speech rate measure from our method for influential ROIs. Given that the $\gamma_{v,p}$ values are assigned a continuous shrinkage prior, we need to apply post-processing on the MCMC samples of $\gamma_{v,p}$ to determine influential voxels following the approach described in the literature (Li and Pati, 2017). While we present inference on $\beta_{v,p}$, we do not emphasize voxel identification in the simulation study following the post-processing steps with $\gamma_{v,p}$ due to space constraint, since our main focus is identifying influential ROIs.
@@ 4 Posterior Computation
Full conditional distributions for all the parameters are available and mostly correspond to standard families (available in the supplementary file). Thus, posterior computation can proceed through a Markov chain Monte Carlo (MCMC) algorithm. The MCMC sampler is run for $5000$ iterations, with the first $1000$ discarded as burn-in. All simulation scenarios show an average effective sample size of over 3143, indicating fairly uncorrelated post-burn-in MCMC samples. Different replications of the model are implemented under a parallel architecture by making use of the packages doparallel and foreach within R.
@@
One important point to note is that $\sum_{l=1}^{L}\xi_{p}^{(l)}/L$ empirically approximates $P(\xi_{p}=1|\mbox{Data})$, and offers the uncertainty of identifying the $p$th ROI to be influential. Offering uncertainty in region selection is a salient feature of the proposed model. To determine whether a region is influential, a threshold is required. We have adopted the median probability rule (Barbieri and Berger, 2004) of $t=0.5$, which is widely adopted in the literature. However, we acknowledge that this cut-off may not always be optimal when variables are highly correlated (Barbieri et al., 2021), a consideration that lies beyond the scope of this article. Since there is no theoretically or experimentally validated cut-off specific to our model, this choice was the most natural. Nonetheless, alternative thresholds may be selected based on the scientific context of a given study.
@@ 5.1 Data Generation
For each sparsity level and co-localization degree, the coefficient corresponding to the edge connecting the $p$-th and $p^{\prime}$-th region is drawn from the mixture distribution
$\theta_{(p,p^{\prime}),t}|\xi^{\theta}_{p,t},\xi^{\theta}_{p^{\prime},t}\sim\xi_{p,t}^{\theta}\xi_{p^{\prime},t}^{\theta}N(\mu_{\theta,(p,p^{\prime})},\sigma_{\theta}^{2})+(1-\xi_{p,t}^{\theta}\xi_{p^{\prime},t}^{\theta})\delta_{0},\>\>\>\theta_{(p,p^{\prime}),t}=\theta_{(p^{\prime},p),t}.$ (6)
@@
(6) ensures that any edge connecting to the $p$-th region in the network response is unrelated to the predictor if the $p$-th region is un-influential, i.e., $\xi^{\theta}_{p,t}=0\Rightarrow\theta_{(p,p^{\prime}),t}=0$ for all $p^{\prime}\in\{1,..,P\}$. Similarly, for each un-influential region $\mathcal{R}_{p}$, i.e., when $\xi^{\beta}_{p,t}=0$, the $V$ dimensional GM coefficient ${\boldsymbol{\beta}}_{p,t}$ is set at ${\boldsymbol{0}}$. When $\xi^{\beta}_{p,t}=1$, i.e., the $p$-th region is influential, we randomly choose $\upsilon_{t}$ proportion of cell coefficients in the $p$-th region to be nonzero and the rest are set at zero. Thus $\upsilon_{t}$ is the density of activated voxels within a region. The nonzero coefficients within ${\boldsymbol{\beta}}_{p,t}$ are simulated from $N(\mu_{\beta,(v,p)},\sigma_{\beta}^{2})$, with $\mu_{\theta,(p,p^{\prime})}$ and $\mu_{\beta,(v,p)}$ are drawn from $Unif(0.25,1)$. We fix $\sigma_{\theta}^{2}=1$, $\sigma_{\beta}^{2}=1$. We control the signal-to-noise ratio by changing the error variances $\tau_{t,\theta}^{2}$ and $\tau_{t,\beta}^{2}$.
@@
(A2) Spatial simulations: Simulating true coefficients ${\boldsymbol{\Theta}}_{t}$ and ${\boldsymbol{\beta}}_{p,t}$. We also consider simulation scenarios where structural coefficients are spatially indexed, i.e., $\beta_{v,p}=\beta({\boldsymbol{s}}_{v,p})$, with ${\boldsymbol{s}}_{v,p}$ representing the spatial location of the $v$th voxel in the $\mathcal{R}_{p}$. This simulation setting assumes perfect co-localization, i.e., $\xi_{p,t}=\xi_{p,t}^{\theta}=\xi_{p,t}^{\beta}$. The network edge coefficient $\theta_{(p,p^{\prime}),t}$ is simulated following (6). When $\xi_{p,t}=1$, we randomly choose $\upsilon_{t}$ proportion of voxel coefficients in $\mathcal{R}_{p}$ to be nonzero, and the rest are set at zero. the nonzero coefficient $\beta({\boldsymbol{s}}_{v,p})$ is simulated from a Gaussian process with mean 1 and a covariance function given by the Matérn Kernel: $cor(\beta({\boldsymbol{s}}_{v,p}),\beta({\boldsymbol{s}}_{v^{\prime},p}))=K(d|\rho,\eta)=\frac{2^{1-\eta}}{\Gamma(\eta)}\left((2\eta)^{1/2}\frac{d}{\rho}\right)^{\eta}K_{\eta}\left((2\eta)^{1/2}\frac{d}{\rho}\right),$ where $d=||{\boldsymbol{s}}_{v,p}-{\boldsymbol{s}}_{v^{\prime},p}||$ and $K_{\eta}$ is the modified Bessel function of order $\eta$. We set $\eta=2$ to ensure a fairly smooth spatial surface and $\rho=1$ for a reasonable range of spatial dependence.
@@
We simulate data under different node sparsity parameters $(1-\nu_{t})$, different signal-to noise ratio, different degree of co-localization $c_{t}$ and different density of activated voxels $\upsilon_{t}$ within an ROI. Finally, a different number of ROIs and different numbers of voxels within an ROI are considered. Specifically, we consider two cases, (a) $P=20$ and $V=10$, and (b) $P=100$ and $V=50$. Cases (a) and (b) are referred to as the “small dimensional example” and “high dimensional example.” All simulation scenarios are summarized in Table 1.